OpenStudy (ell):

Medal to a helper! Please! Name the property illustrated below: a. distributive property b. commutative property c. associative property d. additive inverse e. multiplicative inverse 1. 7+(-3)=(-3)+7 2. 8(-2+12)=-16+96 3. 1/4 * 4 = 1 4. 2* (3*5)= (2*3)*5

OpenStudy (anonymous):

2. distributive property.. because you're distributing 8 to the -2 and then +12 1. Commutative property 4. Associative property 3. Multiplicative inverse

OpenStudy (anonymous):

1. Commutative Property 2. Distributive Property 3. Multiplicative Property 4. Associative Property.

OpenStudy (ell):

wouldn't you distribute the 8?

OpenStudy (ell):

@mathmath333 could you please help me to understand?

OpenStudy (anonymous):

Yes you distribute the 8

OpenStudy (mathmath333):

1. 7+(-3)=(-3)+7 it matches with commutative a+b=b+a

OpenStudy (mathmath333):

8(-2+12)=-16+96 8(-2+12)=-2*8+12*8 it mathches with distributive a(b+c)=ab+ac

OpenStudy (ell):

I've got it guys. Didnt realize how easy it is. Thanks! :) Godbless!